# ConFit v2: Improving Resume-Job Matching using Hypothetical Resume Embedding and Runner-Up Hard-Negative Mining

### v2

@@ Abstract
Xiao Yuthanks: denotes equal contribution  Ruize Xu∗♠  Chengyuan $\vardiamondsuitabsent\vardiamondsuit{}^{*\vardiamondsuit}$  Jinzhong Zhang  Zhou Yu Columbia University  $\vardiamondsuit\vardiamondsuit{}^{\vardiamondsuit}$niversity of Toronto  Intellipro Group Inc. {xy2437,zy2461}@columbia.edu {jinzhong}@intelliprogroup.com
@@ 1 Introduction
Figure 1: Performance comparison between ConFit v2 and ConFit across the AliYun and IntelliPro dataset. “Rank $R$” indicates ranking resume, and “Rank $J$” indicates ranking job.
@@ 2 Background
Despite the rapid growth of online recruitment platforms and the increasing availability of job posts and resume data, there are very few interaction labels between any resumes and job posts Bian et al. (2020); Yu et al. (2024b). This is because a candidate usually applies to only a few positions, and a job interviewer often reviews only a few resumes for a given job post. Often, the resulting dataset $\mathcal{D}=\{R_{i},J_{i},y_{i}\}$ has size $|\mathcal{D}|\ll n_{R}\times n_{J}$, where $n_{R}$ and $n_{J}$ are the total number of resumes and jobs respectively, and $y_{i}\in\{0,1\}$ is a binary signal representing whether a resume $R_{i}$ is accepted for an interview by a job $J_{i}$. For example, in both the IntelliPro and AliYun datasets (Section 4.1) used in this work, less than 0.05% of the total possible (resume, job) pairs are annotated. This label sparsity poses challenges in: 1) crafting high-quality training data/hard negatives for training a neural encoder $f_{\theta}$; and 2) learning a representation space that generalizes well across diverse resumes and job posts.
@@ 3.4 ConFit v2
Intellipro Dataset AliYun Dataset
@@
Intellipro Dataset AliYun Dataset
@@ 4.1 Dataset and Preprocessing
Intellipro Dataset
The resumes and job posts are collected from a global hiring solution company, “Intellipro Group Inc.” To protect the privacy of candidates, all records have been anonymized by removing sensitive identity information (see Appendix D). For each resume-job pair, we record whether the candidate is accepted ($y=1$) or rejected ($y=0$) for an interview. Similar to the AliYun dataset, all resumes and jobs were available as a collection of sections/fields. Both English and Chinese resumes and jobs are included.
@@ 4.3 Baselines
Intellipro Dataset AliYun Dataset
@@ 4.4 Main Results
Table 1 summarizes ConFit v2’s performance in comparison to other baselines on the Intellipro and AliYun datasets when E5-base is used as the encoder. We report two ranking metrics, Recall@K and nDCG@K. In general, ConFit shows substantial improvement over many prior methods, such as MV-CoN, InEXIT, and BM25. ConFit v2 further improves ConFit by 17.1% in recall and 20.4% in nDCG score on average, outperforming all other baselines, including OpenAI’s text-embedding-003-large, on both datasets.
@@ 4.5 Iterative RUM
Since RUM only requires an encoder model to mine hard negatives, we also experiment with using RUM iteratively to improve the model’s performance. For each iteration, we 1) use the encoder trained from the previous iteration to mine hard negatives using RUM; and 2) train the backbone encoder again using the newly acquired hard negatives444 We also experimented with continuously training the model from the previous iteration, but found that it can more easily lead to overfitting. . Throughout all iterations, we keep all hyperparameters the same as in Section 4.2. We focus on the ranking resume task in the IntelliPro dataset since it is the most challenging.
@@ 5 Analysis
In this section, we provide both qualitative and quantitative analysis on the ranking output produced by ConFit v2. We mainly focus on the IntelliPro dataset as it is more challenging.
@@ 5.2 Bias Analysis
ConFit v2 relies on pretrained encoders such as E5 and Jina-v2 Devlin et al. (2019); Wang et al. (2022), and it is well-known that many powerful encoders can contain biases Brunet et al. (2019); May et al. (2019); Jentzsch and Turan (2022); Caliskan et al. (2022) such as gender, age, demographics, etc. As a case study, we examine whether ranking outputs from ConFit v2 contain gender bias. For this experiment, we add gender information back to our desensitized resumes, and compare 1) whether training on gender-included data increases gender disparities in the ranking output; and 2) whether RUM has any effect on gender bias.
@@ B.2 HyRe Upperbound
We experiment with three strategies of selecting “ideal” reference resumes for each job from the dataset. For max-match (human), we select the highest-scoring resume for each job according to additional annotations provided by IntelliPro. For max-match (Jina), we select the highest-scoring resume according to the cosine similarity of the resume/job embeddings, computed directly by Jina-v2 Günther et al. (2023). For centroid, we select the centroid resume of all accepted resumes given a job post. Since computing the centroid resume across the full resume pool is time-consuming, we instead restrict the pool to all labeled resumes given the job post. Then, we pick the resume, which has the shortest average $L_{2}$ normalized distance (i.e. cosine similarity) to its job’s accepted resumes, as the centroid resume.
@@ B.4 Finetuning HyRe
In this section, we examine whether fine-tuned LLMs are capable of generating hypothetical resume to approxiate Section B.2. We perform Supervised Fine-tuning (SFT) and Direct Preference Optimization (DPO) on Qwen-2.5 series models Team (2024). Specifically, we train Qwen-2.5-32B-Instruct with LoRA Hu et al. (2022) on IntelliPro Dataset. All the experiments are implemented on 4 A100 80GB GPUs.
@@
To speed up evaluation, we follow Section B.2 and evaluate all methods using a subset of the IntelliPro dataset. We present the result in Table A3. We find after SFT+DPO training. the Qwen model improves ConFit, but it does not overpass Few-shot prompting on GPT-4o-mini. We believe this is because accepting a resume is subjective, making it hard to learn an “ideal candidate” directly from the dataset itself.
Intellipro Dataset
@@
Train Intellipro Dataset Aliyun Dataset
@@
Intellipro Dataset AliYun Dataset
@@ Appendix D Details on Dataset and Preprocessing
Intellipro Dataset
The talent-job pairs come from the headhunting business in Intellipro Group Inc. The original resumes/job posts are parsed into text fields using techniques such as OCR. Some of the information is further corrected by humans. All sensitive information, such as names, contacts, college names, and company names, has been either removed or converted into numeric IDs. Example resume and job post are shown in Table A9 and Table A10, respectively.
@@
The 2019 Alibaba job-resume intelligent matching competition provided resume-job data that is already desensitized and parsed into a collection of text fields. There are 12 fields in a resume (Table A9) and 11 fields in a job post (Table A10) used during training/validation/testing. Sensitive fields such as “{CJK}UTF8gbsn居住城市” (living city) were already converted into numeric IDs. “{CJK}UTF8gbsn工作经验” (work experience) was processed into a list of keywords. Overall, the average length of a resume or a job post in the AliYun dataset is much shorter than that of the Intellipro dataset (see Table A5).
@@
$R$ from Intellipro Dataset $R$ from AliYun Dataset
@@
Table A9: Example resume from the Intellipro dataset and AliYun dataset. The Intellipro dataset contains resumes in both English and Chinese, while the AliYun dataset contains resumes only in Chinese. All documents are available as a collection of text fields, and are formatted into a single string as shown above for ConFit training.
$J$ from Intellipro Dataset $J$ from AliYun Dataset
@@
Table A10: Example job post from the Intellipro dataset and AliYun dataset. The Intellipro dataset contains job posts in both English and Chinese, while the AliYun dataset contains job posts only in Chinese. All documents are available as a collection of text fields, and are formatted into a single string as shown above for ConFit training.

### v3

@@ Abstract
Xiao Yuthanks: denotes equal contribution Ruize Xu∗♠ Chengyuan $\vardiamondsuitabsent\vardiamondsuit{}^{*\vardiamondsuit}$ Jinzhong Zhang Xu Ma Zhou Yu Columbia University  $\vardiamondsuit\vardiamondsuit{}^{\vardiamondsuit}$niversity of Toronto  Intellipro Group Inc. {xy2437,zy2461}@columbia.edu {jinzhong}@intelliprogroup.com
@@ 1 Introduction
Figure 1: Performance comparison between ConFit v2 and ConFit across the AliYun and the Recruiting dataset. “Rank $R$” indicates ranking resume, and “Rank $J$” indicates ranking job.
@@ 2 Background
Despite the rapid growth of online recruitment platforms and the increasing availability of job posts and resume data, there are very few interaction labels between any resumes and job posts Bian et al. (2020); Yu et al. (2024b). This is because a candidate usually applies to only a few positions, and a job interviewer often reviews only a few resumes for a given job post. Often, the resulting dataset $\mathcal{D}=\{R_{i},J_{i},y_{i}\}$ has size $|\mathcal{D}|\ll n_{R}\times n_{J}$, where $n_{R}$ and $n_{J}$ are the total number of resumes and jobs respectively, and $y_{i}\in\{0,1\}$ is a binary signal representing whether a resume $R_{i}$ is short listed for a job $J_{i}$. For example, in both the Recruiting and AliYun datasets (Section 4.1) used in this work, less than 0.05% of the total possible (resume, job) pairs are annotated. This label sparsity poses challenges in: 1) crafting high-quality training data/hard negatives for training a neural encoder $f_{\theta}$; and 2) learning a representation space that generalizes well across diverse resumes and job posts.
@@ 3.4 ConFit v2
Recruiting Dataset AliYun Dataset
@@
Recruiting Dataset AliYun Dataset
@@ 4.1 Dataset and Preprocessing
Recruiting Dataset
The resumes and job posts are provided by a hiring solution company. To protect the privacy of candidates, all records have been anonymized by removing sensitive identity information (see Appendix D). For each resume-job pair, we record whether the candidate is accepted ($y=1$) or rejected ($y=0$) for an interview. Similar to the AliYun dataset, all resumes and jobs were available as a collection of sections/fields. Both English and Chinese resumes and jobs are included.
@@ 4.3 Baselines
Recruiting Dataset AliYun Dataset
@@ 4.4 Main Results
Table 1 summarizes ConFit v2’s performance in comparison to other baselines on the Recruiting and AliYun datasets when E5-base is used as the encoder. We report two ranking metrics, Recall@K and nDCG@K. In general, ConFit shows substantial improvement over many prior methods, such as MV-CoN, InEXIT, and BM25. ConFit v2 further improves ConFit by 17.1% in recall and 20.4% in nDCG score on average, outperforming all other baselines, including OpenAI’s text-embedding-003-large, on both datasets.
@@ 4.5 Iterative RUM
Since RUM only requires an encoder model to mine hard negatives, we also experiment with using RUM iteratively to improve the model’s performance. For each iteration, we 1) use the encoder trained from the previous iteration to mine hard negatives using RUM; and 2) train the backbone encoder again using the newly acquired hard negatives444 We also experimented with continuously training the model from the previous iteration, but found that it can more easily lead to overfitting. . Throughout all iterations, we keep all hyperparameters the same as in Section 4.2. We focus on the ranking resume task in the Recruiting dataset since it is the most challenging.
@@ 5 Analysis
In this section, we provide both qualitative and quantitative analysis on the ranking output produced by ConFit v2. We mainly focus on the Recruiting dataset as it is more challenging.
@@ 5.2 Bias Analysis
ConFit v2 relies on pretrained encoders such as E5 and Jina-v2 Devlin et al. (2019); Wang et al. (2022), and it is well-known that many powerful encoders can contain biases Brunet et al. (2019); May et al. (2019); Jentzsch and Turan (2022); Caliskan et al. (2022) such as gender, age, demographics, etc. As a case study, we examine whether ranking outputs from ConFit v2 contain gender bias. For this experiment, we add gender information [^5] Since gender information is voluntarily submitted by the candidates, this affects around 15% of all resumes. back to our desensitized resumes, and compare 1) whether training on gender-included data increases gender disparities in the ranking output; and 2) whether RUM has any effect on gender bias.
@@ B.2 HyRe Upperbound
We experiment with three strategies of selecting “ideal” reference resumes for each job from the dataset. For max-match (human), we select the highest-scoring resume for each job according to additional annotations provided by a hiring solution company. For max-match (Jina), we select the highest-scoring resume that achieves high human annotation score and a high cosine similarity score computed directly by Jina-v2 Günther et al. (2023). For centroid, we select the centroid resume of all accepted resumes given a job post. Since computing the centroid resume across the full resume pool is time-consuming, we instead restrict the pool to all labeled resumes given the job post. Then, we pick the resume, which has the shortest average $L_{2}$ normalized distance (i.e. cosine similarity) to its job’s accepted resumes, as the centroid resume.
@@ B.4 Finetuning HyRe
In this section, we examine whether fine-tuned LLMs are capable of generating hypothetical resume to approxiate Section B.2. We perform Supervised Fine-tuning (SFT) and Direct Preference Optimization (DPO) on Qwen-2.5 series models Team (2024). Specifically, we train Qwen-2.5-32B-Instruct with LoRA Hu et al. (2022) on the Recruiting Dataset. All the experiments are implemented on 4 A100 80GB GPUs.
@@
To speed up evaluation, we follow Section B.2 and evaluate all methods using a subset of the Recruiting dataset. We present the result in Table A3. We find after SFT+DPO training. the Qwen model improves ConFit, but it does not overpass Few-shot prompting on GPT-4o-mini. We believe this is because accepting a resume is subjective, making it hard to learn an “ideal candidate” directly from the dataset itself.
Recruiting Dataset
@@
Train Recruiting Dataset Aliyun Dataset
@@
Recruiting Dataset AliYun Dataset
@@ Appendix D Details on Dataset and Preprocessing
Recruiting Dataset
The talent-job pairs are provided by a hiring solution company. The original resumes/job posts are parsed into text fields using techniques such as OCR. Some of the information is further corrected by humans. All sensitive information, such as names, contacts, college names, and company names, has been either removed or converted into numeric IDs. Example resume and job post are shown in Table A9 and Table A10, respectively.
@@
The 2019 Alibaba job-resume intelligent matching competition provided resume-job data that is already desensitized and parsed into a collection of text fields. There are 12 fields in a resume (Table A9) and 11 fields in a job post (Table A10) used during training/validation/testing. Sensitive fields such as “{CJK}UTF8gbsn居住城市” (living city) were already converted into numeric IDs. “{CJK}UTF8gbsn工作经验” (work experience) was processed into a list of keywords. Overall, the average length of a resume or a job post in the AliYun dataset is much shorter than that of the Recruiting dataset (see Table A5).
@@
$R$ from Recruiting Dataset $R$ from AliYun Dataset
@@
Table A9: Example resume from the Recruiting dataset and AliYun dataset. The Recruiting dataset contains resumes in both English and Chinese, while the AliYun dataset contains resumes only in Chinese. All documents are available as a collection of text fields, and are formatted into a single string as shown above for ConFit training.
$J$ from Recruiting Dataset $J$ from AliYun Dataset
@@
Table A10: Example job post from the Recruiting dataset and AliYun dataset. The Recruiting dataset contains job posts in both English and Chinese, while the AliYun dataset contains job posts only in Chinese. All documents are available as a collection of text fields, and are formatted into a single string as shown above for ConFit training.